The electronics company plans to replace the manually operated manufacturing machine with a new fully automated machine. Use the following information to determine the cash flows and profitability of the replacement decision.

Current situation

• Current estimated wages of operators are 30,000 annually. By replacing, we can save these labour related costs.

• Maintenance costs € 8,000 per year.

• Waste related costs of € 10,000 per year

• The machine currently in use was bought 5 years ago for 60,000 euros. The company can continue to operate this machine for the next five years. The company applies linear depreciation and both the book and market values are zero after five years.

• The potential sales price of the old machine today is € 20,000

Project under consideration

  • Required capital investment is € 120,000
  • Maintenance costs € 14,000 per year.
  • Waste related costs of € 5,000 per year
  • The machine is going to be operated for the next 5 years
  • Linear depreciation is applied. Accounting salvage value is zero after 5 years of operation. However, it is estimated that the potential market value of the new machine is 25% of the purchase price.
  • Replacement requires additional investments of € 12,000 into net working capital, which is expected to be recovered when project ends.
  • The company applies 10% cost of capital for this type of replacement projects
  • Corporate tax rate is 40% on both profits and capital gains.

Find:

a) Incremental cash flows from replacing the machine (including investment, annual cash flow, and closing cash flow)

d) Based on the NPV, assess whether the replacement is economically viable
Answer #1

Note -

1. Assuming 25% of 120,000 capital investment is recovered at the end of 5th year.
2. Extra maintenance cost of new project = 14000 - 8000 = 6000
3. Waste Cost saved by new option = 10000 - 5000 = 5000
4. wages saved = 30000
5. Net Capital Required for new machinery = 120,000 - 20,000 = 100,000
6. Working Capital Investment of 12000 is recovered at the end of 5th year
7. Corporate tax rate of 40% is assumed to be already incorporated in the Cost of Capital of 10%.
8. As life of 1st machine was 5 years, no book value is considered for it in calculation for next 5 years.
9. 2nd Machine also has life of 5 years, no book value left at the end of 5 years.
10. Depreciation is a non-cash expense not required for cash flow calculation.
11. NPV is calculated using usual NPV formula.

As, this project has a positive NPV of $27,272.73, project must be taken up.